# An Evaluation of N-Gram Selection Strategies for Regular Expression Indexing in Contemporary Text Analysis Tasks [Experiment, Analysis & Benchmark]

### v1

@@ 7. Learnings and Future Work
The evaluation shows that the optimal choice of n-gram selection strategies among FREE, BEST, and LPMS is strongly influenced by workload characteristics (such as the query set size, query literal sizes, alphabet size, data patterns, etc.). Since no single strategy suits all workloads, this analysis emphasizes the importance of selecting methods based on workload characteristics and provides a general guideline. We summarize the insights in the decision flow chart in Figure 4 as a guideline for practitioners. This guideline recommends 1) choosing FREE for large and diverse workloads or when unseen queries are expected, 2) choosing BEST when precision is crucial for repeated queries and the number of candidate n-grams is small, and 3) choosing LPMS for formatted datasets or when query literals are common in the dataset.

### v2

@@ 7. Learnings and Future Work
The evaluation shows that the optimal choice of n-gram selection strategies among FREE, BEST, and LPMS is strongly influenced by workload characteristics (such as the query set size, query literal sizes, alphabet size, data patterns, etc.). Since no single strategy suits all workloads, this analysis emphasizes the importance of selecting methods based on workload characteristics and provides a general guideline. We summarize the insights in the decision flow chart in Figure 6 as a guideline for practitioners. This guideline recommends 1) choosing FREE for large and diverse workloads or when unseen queries are expected, 2) choosing BEST when precision is crucial for repeated queries and the number of candidate n-grams is small, 3) choosing LPMS for formatted datasets or when query literals are common in the dataset, and 4) choosing VGGraph when precision requirement is high with low index space constraint.
@@ 8. Conclusion
Appendix A FREE Regex Query Plan Compiler
We provide an example index search query plan in Figure 7 for the example regex, <ahref=("|’).*ZZZ\. pdf("|’)>, assuming that only ", , Z, and pdf are indexed. The literal tree shown in Figure 7(a) is built solely by extracting all literals from the regex. Referring to the index, we notice that <ahref= and > are not indexed, and two n-grams in the literal ZZZ.pdf, Z and pdf are indexed. We simplify the plan to get Figure 7(b).
Appendix B Robustness Test
For this test, we use the synthetic workload with different sets of regexes for index construction and query matching to test the robustness of each method for unseen queries.
To test the robustness of the n-gram selection methods in handling unseen queries, we build off the synthetic dataset in the original LPMS (Tsang and Chawla, 2011) paper. The dataset consists of 5000 strings each constructed with alphabet ‘A-P’. The string size follows a geometric distribution with $p=1/32$. The index building query set is generated by randomly selecting 10% from the dataset, and a random slice is used to create a regex lit1regexlit2. lit1 has 1-5 characters, lit2 has 0-5 characters, and regex matches any $m$ characters where $1\leq m\leq 50$. During the query phase, we generate regexes of a similar format from 2% of the data records.
We summarize the result in Appendix B. Overall, FREE demonstrates robustness in this experiment, achieving highest precision with the lowest computational overhead. This can be attributed to the fact that FREE selects n-grams independent of the query set, and thus adapt smoothly to unseen regexes in query time.
For $K=20$, FREE($\max_{n}=2$, $c=0.7$) builds an index with 16 keys that achieves the highest precision of 0.267. It also has the lowest index construction overhead. BEST($c=0.2$) achieves slightly lower precision of 0.208, but the query time is $4.4\times$ more than that of FREE. LPMS-D performs the worst with the lowest precision and highest query time. Note that BEST and LPMS also have significantly higher computational overhead compared to FREE.
As $K$ increases to 100, BEST($c=0.5$) achieves the highest precision. It’s index construction time also increases by $\approx$3 times to 31.7s. This is more than $1000\times$ longer than FREE. Note that the query times for BEST are not the lowest among the three methods, although it has the highest precision. This is due to the non-uniformness in the data record lengths, which breaks Assumption Assumption BEST-2.
As we increase to $K=300$, FREE($max_{n}=2$, $c=0.12$) achieves the highest precision of 0.6453. BEST($c=0.2$) achieves similar precision and similar overall query time compared to FREE, but with a much larger index construction time which is $2833\times$ slower than FREE. In all cases, LPMS performs unfavorable to the other two methods, since the n-gram selection strategy of LPMS is solely based on the benefit calculation using the query set and the dataset, without any heuristic like the pruning step of BEST.
{customframe}
Insights
Table 7. Index cost and query performance on Robust.
$K$ Method $T_{Q}$ s $T_{I}$ s $S_{Q}$GB $S_{I}$MB PrecPrec
20 BEST 0.829 3.398 0.042 0.091 0.2084
[…]